OpenStudy (anonymous):

find the midpoint for the segment shown (-3,-5), (1,8)

OpenStudy (anonymous):

Okay. It's simply to find the middle between those two. So let's do 'x' first. Between -3 and 1 on a table is 4. [4 is called the difference], so we half 4 and add it on to the lower, to find the midpoint of x. Half of four is two. So -3 + 2 = -1. -1 is the x midpoint. Now let's do 'y'. Between -5 and 8 is a difference of '13'. So we have it, and add it on to the lower one. Half of 13 is 6.5. Add 6.5 onto -5 and you get 1.5. So, the coordinate for the midpoint is; (-1,1.5)

OpenStudy (anonymous):

Use the formula M = (x1 + x2 / 2 , y1 + y2 / 2) M = ( -3 + 1 / 2 , -5 + 8 /2) You'll end up with M = ( -2/2 , 3/2) which is ( -1, 1.5) Therefore, the midpoint lies on ( -1, 1.5).